Loan Interest Rates: A realtor borrowed $90,000 to develop some property. He was able to borrow part of the money at 5.5% interest and the rest at 6%. The annual interest on the two loans amounts to $5125. How much was borrowed at each rate?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

0.055x+0.06(90,000-x) = 5125
0.055x+5,400-0.06x = 5125
0.005x = 275
x = 275/0.005 = 55,000
Borrowed:
55,000 at 5.5%
35,000 at 6%

Answer by ikleyn(52776)

.
To see solved problems/samples of this type, look into the lesson
    - Using systems of equations to solve problems on investment
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Systems of two linear equations in two unknowns".